OpenStudy (anonymous):

How do I get the answer for 1/2+(2/3 / 3/4) - (4/5 * 5/6) = The way I've worked it is Divide first so 2/3 / 3/4 introverted and multiplied is 8/9 then 4/5*5/6 is 20/30 so now I have 1/2 + 8/9 - 20/30 what do I do now?

OpenStudy (anonymous):

you have 1/2 + 8/9 - 20/30 simplified gives 1/2 + 8/9 - 2/3 you need to find a common denominator, which would be 18 so you now would have 9/18+16/18-12/18 now you just add the numerators so the answer is (9+18-12)/18 = 15/18 = 5/6

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

nws